OpenStudy (anonymous):

In everyday experience, the measures of temperature most often used are Fahrenheit F and Celsius C. Recall that the relationship between them is given by the following formula. F = 1.8C + 32 Physicists and chemists often use the Kelvin temperature scale. You can get kelvins K from degrees Celsius by using the following formula. K = C + 273.15 1. Find a formula expressing the temperature F in degrees Fahrenheit as a function of the temperature K in kelvins.

OpenStudy (skullpatrol):

Any ideas :-)

OpenStudy (skullpatrol):

Since the question asks to find a formula expressing the temperature F in degrees Fahrenheit as a function of the temperature K in kelvins you need to eliminate the temperature measured in Celsius C.

OpenStudy (skullpatrol):

These are the two given formulas: $$\Huge F = 1.8C + 32$$ $$\Huge K = C + 273.15$$

OpenStudy (skullpatrol):

Note the question asks for F as a function of K which means F(K).

OpenStudy (anonymous):

Right..

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

Not at all, you said we had to eliminate the C but I have no idea how lol

OpenStudy (skullpatrol):

Choose one of the equations and solve it for C.

OpenStudy (skullpatrol):

Can you do that?

OpenStudy (anonymous):

So, F-32/1.8 = C

OpenStudy (skullpatrol):

Ok, now substitute that into the other equation and simplify.

OpenStudy (anonymous):

THANK YOU SO MUCH OMG

OpenStudy (skullpatrol):